# Single-beam grating-chip 3D and 1D optical lattices

### v1

@@ Abstract
Fig. 1 a illustrates the central of the GMOT chip we use to create a 3D lattice. The chip comprises three sectors of one-dimensional diffraction grating (gray triangles). The binary grating sectors have properties and fabrication methods as detailed in [36, 45, 46, 41], and are optimised for the atomic species utilised here – 87Rb. We use a similar theoretical description to our MOT modelling papers [35, 47], however here we consider the coherent superposition of electric fields rather than the Doppler- and Zeeman-effect sensitive sums of intensity-based average radiation pressures (see Supplementary Material (SM)). Theoretical cross-sections of the 3D lattice itself are shown in Fig. 1 a and b, illustrating the different lattice regions with their distinct geometries and dimensions. The 1D lattice model is shown in Fig. 1 c, with experimental data in d.
@@
As a basis for loading the optical lattices, 8 87Rb atoms are loaded into a GMOT and cooled to 3K using a standard red molasses cooling stage [36, 37]. Stray DC magnetic fields are cancelled by three pairs of orthogonal Helmholtz coils. The MOT location with respect to the chip, and its overlap with the optical lattice, can be controlled within the GMOT’s optical overlap volume by varying the zero position of the quadrupole magnetic field. The typical MOT location and size are also shown in Fig. 1 a and b.
@@
Atom detection is achieved via fluorescence imaging onto a CCD camera sensor. The imaging system is aligned orthogonal to the directions of both the MOT/lattice beam and gravity. A spatially selective focal plane is utilised to reduce background noise originating from scattered light 333Additionally, a notch filter with a peak transmission centred around nm is used when investigating the 1D nm lattice. For both the 3D and 1D lattice geometries s long resonant fluorescent images are taken with the MOT light after a ms time of flight following the extinction of the lattice beam. A calculation of atomic motion during imaging (see SM) quantitatively matches the large measured width of the cloud seen in Fig. 1 d. In both the 3D and 1D lattices, the lattice beam power can be rapidly adjusted by a single-pass acousto-optic modulator (AOM) designed for the beam wavelength.
@@
A log plot of the number of trapped atoms as a function of hold time in both the 3D and 1D lattice is shown in Fig. 1 e. Exponential decays are fitted, excluding data where ms, to eliminate untrapped atoms ejected from the imaging region over this time scale due to gravity. The 1D lattice has a much longer lifetime, this is because it is much deeper and the far-off resonance light also induces less scattering and thereby heating than the near-resonant 3D lattice. The overlap volume between the 3D lattice and the MOT is much larger than for the 1D lattice, which can be seen by the difference in initial lattice atom number.
@@
The measured parametric resonance frequencies in the 1D lattice (Fig. 2 c) are kHz and kHz, which, like in the 3D case, match well with sub-harmonics of the expected trap frequencies, kHz and kHz. There should be much less spatial variation in trap frequency due to lattice structure in the 1D lattice compared to the 3D lattice, and the atoms are expected to be in the deeper harmonic parts of the lattice, with no trap frequency variation with state. In the 1D case, however, the signal-to-noise of the parametric heating is low, and the (relatively broad) resonances can only be observed via atom number loss. This is due to lower spatial overlap of the 1D lattice with the MOT and hence lower initial atom number (Fig. 1 e), as well as rapid radial expansion during fluorescence imaging (Fig. 1 c, d and the SM). The 1D lattice beam focal position on the grating surface was optimised for atom number, and given the lattice beam Rayleigh range of mm, the beam waist was constant over the entire lattice.
In this paper we have demonstrated optical lattices in one and three dimensions. However, we note that by loading the MOT atoms into a magnetic and/or optical trap, they can be transported beyond the MOT capture volume, and thus future single-input-beam two-dimensional lattices are within easy reach (Fig. 1 a, b).
@@ Acknowledgements.
The authors would like to thank Arthur La Rooij and Oliver Burrow for valuable discussions. A.B. was supported by a Ph.D. studentship from the Defence Science and Technology Laboratory (Dstl). J.P.M. gratefully acknowledges support from a Royal Academy of Engineering Research Fellowship. We thank the UK EPSRC for funding via grant EP/T001046/1. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript (AAM) version arising from this submission. The dataset is available here 777TBD.References
@@
- Note [7] TBD.

### v2

@@ Abstract
Fig. 1 a illustrates the central of the GMOT chip we use to create a 3D lattice. The chip comprises three sectors of one-dimensional diffraction grating (gray triangles). The binary grating sectors have properties and fabrication methods as detailed in [36, 45, 46, 41], and are optimised for the atomic species utilised here – 87Rb. We use a similar theoretical description to our MOT modelling papers [35, 47], however here we consider the coherent superposition of electric fields rather than the Doppler- and Zeeman-effect sensitive sums of intensity-based average radiation pressures (see Supplementary Material (SM)). Theoretical cross-sections of the 3D lattice itself are shown in Fig. 1 b and c, illustrating the different lattice regions with their distinct geometries and dimensions. The 1D lattice model and experimental data are shown in Fig. SM2 a and b, respectively.
@@
As a basis for loading the optical lattices, 8 87Rb atoms are loaded into a GMOT and cooled to 3K using a standard red molasses cooling stage [36, 37]. Stray DC magnetic fields are cancelled by three pairs of orthogonal Helmholtz coils. The MOT location with respect to the chip, and its overlap with the optical lattice, can be controlled within the GMOT’s optical overlap volume by varying the zero position of the quadrupole magnetic field. The typical MOT location and size are also shown in Fig. 1 b and c.
@@
Atom detection is achieved via fluorescence imaging onto a CCD camera sensor. The imaging system is aligned orthogonal to the directions of both the MOT/lattice beam and gravity. A spatially selective focal plane is utilised to reduce background noise originating from scattered light 333Additionally, a notch filter with a peak transmission centred around nm is used when investigating the 1D nm lattice. For both the 3D and 1D lattice geometries s long resonant fluorescent images are taken with the MOT light after a ms time of flight following the extinction of the lattice beam. In both the 3D and 1D lattices, the lattice beam power can be rapidly adjusted by a single-pass acousto-optic modulator (AOM) designed for the beam wavelength.
@@
A log plot of the number of trapped atoms as a function of hold time in both the 3D and 1D lattice is shown in Fig. 2. Exponential decays are fitted, excluding data where ms, to eliminate untrapped atoms ejected from the imaging region over this time scale due to gravity. The 1D lattice has a much longer lifetime, this is because it is much deeper and the far-off resonance light also induces less scattering and thereby heating than the near-resonant 3D lattice. The overlap volume between the 3D lattice and the MOT is much larger than for the 1D lattice, which can be seen by the difference in initial lattice atom number.
@@
The measured parametric resonance frequencies in the 1D lattice (Fig. 3 c) are kHz and kHz, which, like in the 3D case, match well with sub-harmonics of the expected trap frequencies, kHz and kHz. There should be much less spatial variation in trap frequency due to lattice structure in the 1D lattice compared to the 3D lattice, and the atoms are expected to be in the deeper harmonic parts of the lattice, with no trap frequency variation with state. In the 1D case, however, the signal-to-noise of the parametric heating is low, and the (relatively broad) resonances can only be observed via atom number loss. This is due to lower spatial overlap of the 1D lattice with the MOT and hence lower initial atom number (Fig. 2), as well as rapid radial expansion during fluorescence imaging (Fig. SM2 and discussion in the SM). The 1D lattice beam focal position on the grating surface was optimised for atom number, and given the lattice beam Rayleigh range of mm, the beam waist was constant over the entire lattice.
In this paper we have demonstrated optical lattices in one and three dimensions. However, we note that by loading the MOT atoms into a magnetic and/or optical trap, they can be transported beyond the MOT capture volume, and thus future single-input-beam two-dimensional lattices are within easy reach (Fig. 1 b, c).
@@ Acknowledgements.
The authors would like to thank Arthur La Rooij and Oliver Burrow for valuable discussions. A.B. was supported by a Ph.D. studentship from the Defence Science and Technology Laboratory (Dstl). J.P.M. gratefully acknowledges support from a Royal Academy of Engineering Research Fellowship. We thank the UK EPSRC for funding via grant EP/T001046/1. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript (AAM) version arising from this submission. The dataset is available here 777Dataset DOI to be added., and supplementary material which includes Refs. [70, 71, 72] can be found here 888Supplementary Material link to be added.References
@@
- Note [7] Dataset DOI to be added.
- Vangeleyn et al. [2010] M. Vangeleyn, P. F. Griffin, E. Riis, and A. S. Arnold, Laser cooling with a single laser beam and a planar diffractor, Optics Letters 35, 3453 (2010).
- Henderson et al. [2020] V. A. Henderson, M. Y. H. Johnson, Y. B. Kale, P. F. Griffin, E. Riis, and A. S. Arnold, Optical characterisation of micro-fabricated Fresnel zone plates for atomic waveguides, Optics Express 28, 9072 (2020).
- Arnold and Manson [2000] A. S. Arnold and P. J. Manson, Atomic density and temperature distributions in magneto-optical traps, Journal of the Optical Society of America B 17, 497 (2000).
- Note [8] Supplementary Material link to be added.